# Conceptualization of this study, Methodology, Software

### v1

@@ 3.2 Muti-Scale Retention
In this study, we explore the integration of the Retentive Network model, originally designed for Natural Language Processing, into the realm of EEG signal denoising. It is important to emphasize that while the Retentive Network has demonstrated remarkable success in natural language processing, our investigation centers on its applicability to EEG denoising. The Retentive Network (Retnet) is comprised of L identical modules arranged in a stacked fashion, featuring residual connectivity and pre-LayerNorm akin to the Transformer architecture. Each Retnet module comprises two sub-modules: the Multi-scale Retention (MSR) module and the Feedforward Network (FFN) module. For a given input sequence $s=s_{1}s_{2}s_{3}\dots s_{\left|s\right|}$ (where $\left|s\right|$ denotes the length of the sequence), the input vector is initially transformed to $X^{0}=\left[x_{1},x_{2},\dots,x_{|x|}\right]\in\mathbb{R}^{|x|\times d_{model}}$, where $d_{model}$ is the hidden dimension. Subsequently, the Retnet Block can be computed for each layer, denoted as $X^{l}=Retnet(X^{l-1}),l\in[1,L]$. The Simple Retention layer is defined as follows [19]:
@@
where $\bar{\Theta}$ is the complex conjugate of $\Theta$, $W_{Q},W_{K},W_{V}\in\mathbb{R}^{|x|\times|x|}$, $D\in\mathbb{R}^{|x|\times|x|}$ combines causal masking and exponential decay of relative distances into one matrix.
To achieve a multichannel-like effect, input sequences can be projected to lower dimensions $d$ times, akin to the multiple-header mechanism in Transformer. This method is employed in each Retention layer with multiple headers $h=\frac{d_{model}}{d}$, where $d$ represents the length of the sequences in each header. Each header utilizes distinct $W_{Q},W_{K},W_{V}\in\mathbb{R}^{|d|\times|d|}$, constituting the Muti-Scale Retention (MSR) block. Different $\gamma$ hyperparameters are assigned to various heads in MSR, maintaining simplicity with the same $\gamma$ across different layers. Additionally, a swish [9] gate is incorporated to enhance nonlinear features in each layer. Given an input $X$, the mathematical representation of the Muti-Scale Retention is provided as follows:
@@
Here, $W_{G},W_{O}\in\mathbb{R}^{d_{model}\times d_{model}}$ are learnable parameters, and GroupNorm normalizes the output of each head.
  • Group Normalization
    Group Normalization (GN) [24] is an alternative to traditional Batch Normalization (BN) [8] that addresses the issue of poor performance with small batch sizes. When using small batch sizes, BN may not effectively estimate the mean and variance of the entire data distribution. Additionally, BN may yield suboptimal results for tasks with highly imbalanced binary classification. To overcome these limitations, GN divides the channels into groups and performs normalization within each group. Instead of computing the mean and variance based on the batch size, GN computes the mean over $(\frac{C}{G})\times H\times W$, where C is the number of channels, G is the number of groups, H is the height, and W is the width. By doing so, GN decouples the normalization from the batch size and allows the network to achieve better performance. By normalizing within groups, GN provides a more robust estimation of statistics and helps mitigate the negative impact of small batch sizes or imbalanced data distribution.
@@ 3.3 Signal Embedding
In our extended investigation, it was observed that when the input sequence $s=s_{1}s_{2}s_{3}\dots s_{|s|}\in\mathbb{R}^{|s|\times 1}$ is relatively short, direct embedding is feasible. However, in general scenarios, where the time-series information of the signal is usually lengthy, direct embedding incurs high computational complexity, hindering effective network training. To address this, we propose the introduction of a concept termed "patch", involving the amalgamation of a series of consecutive samples into a single input feature. This concept is inspired by speech signal processing, where a solitary sample may inadequately represent the current word, while a segment of samples offers more semantic expressiveness. It is noteworthy that EEG signals frequently encompass extensive temporal information, and signal embedding intelligently captures this temporal data. By grouping consecutive samples into patches, the network better retains context and temporal relationships in the signal, enhancing denoising effectiveness. This approach aligns with speech signal processing, where context is pivotal for accurate speech comprehension. Consequently, this paper introduces signal embedding, a more efficient process tailored to the characteristics of EEG signals.
The complete signal embedding process is illustrated in Figure 2(c). Assuming a given patch size, the original sequence is divided accordingly, reducing the sequence length to $|s|//patch\_size$. Subsequently, each patch undergoes reshaping and linear projection to attain the desired hidden dimension. This process not only mitigates computational complexity but also preserves timing information more effectively. It’s crucial to note that the signal embedding used here does not employ positional encoding. This is because the Retention mechanism already incorporates positional encoding considerations, obviating the need for additional positional encoding. Mathematically, it can be expressed as follows:
${s}^{\prime}=Pathchfiy(s)={s}^{\prime}_{1}{s}^{\prime}_{2}{s}^{\prime}_{3}\dots s^{\prime}_{\frac{|s|}{p}}\in\mathbb{R}^{\frac{|s|}{p}\times p}$ (8)
$X^{0}=Embedding({s}^{\prime};\omega)=[x_{1},x_{2},\dots,x_{|x|}]\in\mathbb{R}^{\frac{|s|}{p}\times d_{model}}$ (9)
@@ 4.1 Preliminary
Signals disturbed by noise are acquired through the linear combination of the electrooculogram (EOG) or electromyogram (EMG) with the pristine electroencephalogram (EEG). This procedure can be mathematically represented as Equation (11) [27]. The mixed EEG noise signal is denoted as $y\in\mathbb{R}^{|y|}$, where $y$ represents the sequence length. The noise-free EEG signal, denoted as $x\in\mathbb{R}^{|x|}$, serves as the ground truth, and $n\in\mathbb{R}^{|n|}$ represents oculomotor or electromyographic artifacts (myogenic artifacts). It is important to note that the lengths of each sequence $|x|,|y|,|n|$ are equal. To control the noise level during mixing, we introduce the hyperparameter $\lambda$, regulating the signal-to-noise ratio (SNR) of the noisy signal. Adjustment of different $\lambda$ values enables effective control of SNR magnitude to adapt to various noise environments. The SNR is calculated using Equation (12), while $\lambda$ is determined by Equation (13), where $RMS(\cdot)$ denotes the root mean square of the sample, $RMS(x)$ is the root mean square of the noiseless EEG signal $x$, and RMS $RMS(\lambda\cdot n)$ is the root mean square of the mixed noise $\lambda\cdot n$. These formulas provide flexible adjustment of the signal-to-noise balance to meet diverse signal quality requirements in specific application scenarios.
@@ 4.2.1 Datasets
However, it is important to note that the EEGDenoiseNet dataset solely provides raw data, necessitating researchers to conduct their own processing. This procedure is relatively intricate, posing inconvenience for the exploration of EEG signal denoising through deep learning methodologies. To address this challenge, this paper undertakes the preprocessing of the dataset and subsequently shares the processed dataset as an open-source resource, now accessible on the Hugging Face Hub. The primary objective of this endeavor is to facilitate researchers’ access to and utilization of processed data, enabling them to concentrate more on the investigation of EEG deep learning denoising methods without being encumbered by the intricacies of data processing. By providing this dataset as an open-source entity, our aim is to stimulate increased research in EEG signal processing and offer a more streamlined resource for the academic community.
@@ 4.2.2 Train details
Throughout the training phase, the network underwent 5000 epochs to ensure comprehensive learning of dataset features. Furthermore, we configured the batch size to 1000, a standard choice that balances memory utilization and training efficacy. Notably, for accelerated training, we harnessed the computational capabilities of an NVIDIA GeForce RTX 3090 Graphics Processing Unit (GPU). Leveraging the parallel computing power of GPUs significantly augmented the speed of deep learning model training, facilitating rapid experimentation and tuning for researchers.
@@ 4.3.2 Evaluation measures
We assess the denoising outcomes using three methods: Relative Root Mean Squared Error in the temporal domain ($RRMSE_{temporal}$), RRMSE in the spectral domain ($RRMSE_{spectral}$), and the correlation coefficient ($CC$) [27]. This selection is grounded in a profound understanding of EEG signal characteristics. Firstly, considering the temporal significance of the EEG signal, we employ $RRMSE_{temporal}$ to quantify the relative error between the denoised and original signals in the time domain. This method exhibits sensitivity to denoising techniques preserving temporal information. Secondly, as EEG signals encapsulate rich spectral information with research often focusing on specific frequency ranges, $RRMSE_{spectral}$ is employed to ensure the preservation of features in the frequency domain. Lastly, acknowledging the synergistic activities between different brain regions in EEG signals, $CC$ serves as an evaluation metric. $CC$ reflects the linear relationship between the denoised and original signals, crucial for maintaining vital information about interregional correlation. The combined use of these methods facilitates a comprehensive evaluation of denoising performance across time and frequency domains, considering their adaptability to EEG signal characteristics. The mathematical expressions for $RRMSE_{temporal}$, $RRMSE_{spectral}$, and $CC$ are represented in Equation (16), (17), and (18) respectively, where $RMS(\cdot)$ denotes root mean square, $PSD(\cdot)$ denotes power spectral density, and $Cov(\cdot)$, $Var(\cdot)$ represent covariance and variance, respectively.
$RRMSE_{temporal}=\frac{RMS(F(y)-x)}{RMS(x)}=\frac{RMS(\hat{x}-x)}{RMS(x)}$ (16)
$RRMSE_{spectral}=\frac{RMS(PSD(F(y))-PSD(x))}{RMS(PSD(x))}=\frac{Cov(\hat{x},x)}{Var(\hat{x})Var(x)}$ (17)
@@ 4.3.3 Ablation study
In this study, we pioneer the application of the Retnet network to 1D signal denoising and introduce the innovative concept of signal embedding. To assess the impact of each hyperparameter on denoising performance, we conduct ablation experiments, marking the inaugural exploration of these techniques. Our initial focus is on the influence of patch size and hidden dimension, investigating their effects on network performance. Table 1 presents quantized results for the network applied to EOG and EMG under various configurations of patch size and hidden dimension hyperparameters. Notably, we observe an enhancement in denoising performance with decreasing patch size while maintaining a consistent hidden dimension. This improvement is attributed to the impact of patched sequence length on the network’s feature extraction capability—smaller patch sizes result in larger mini sequence lengths, preserving more information and thereby improving denoising effectiveness. However, a cautious approach is essential as blindly reducing patch size escalates computational complexity due to increased sequence length. Thus, a delicate balance between denoising performance and computational efficiency is imperative. Furthermore, maintaining the same patch size, an increased hidden dimension corresponds to improved denoising performance, aligning with the intuitive understanding that higher dimensionality facilitates enhanced feature extraction.
Table 1: The effect of patch size and hidden dim on the noise reduction performance of EEGDiR. Note mini sequence length must be $|s|//patchsize$, where $|s|=512$, with $8$ Heads and $4$ DiRBlock layers.
Patch size Mini sequence length Hidden dim Ocular artifact Muscle artifact
$RRMSE_{temporal}$ $RRMSE_{spectral}$ $CC$ $RRMSE_{temporal}$ $RRMSE_{spectral}$ $CC$
32 16 512 0.339 0.367 0.928 0.556 0.561 0.793
32 16 256 0.353 0.377 0.911 0.569 0.575 0.791
32 16 64 0.382 0.371 0.909 0.572 0.577 0.790
16 32 512 0.327 0.361 0.932 0.532 0.501 0.807
16 32 256 0.348 0.371 0.925 0.598 0.573 0.776
16 32 64 0.374 0.378 0.912 0.654 0.593 0.701
Subsequently, we assess the impact of varying the number of block layers L on network performance. Table 2 displays the denoising quantization results for the model applied to EOG and EMG datasets, with fixed parameters patch size $16$, hidden dimension $512$, and heads $8$. The observations indicate a gradual enhancement in denoising performance with an increasing number of layers. This improvement is ascribed to the benefits of residual connections, whereby a higher number of layers does not lead to overfitting. The increased network depth contributes to superior feature extraction capabilities.
@@
Layers Ocular artifact Muscle Artifact
$RRMSE_{temporal}$ $RRMSE_{spectral}$ $CC$ $RRMSE_{temporal}$ $RRMSE_{spectral}$ $CC$
4 0.327 0.361 0.932 0.532 0.501 0.807
3 0.356 0.380 0.925 0.578 0.583 0.789
2 0.372 0.383 0.917 0.591 0.596 0.781
1 0.394 0.429 0.908 0.613 0.594 0.766

### v2

@@ 3.2 Muti-Scale Retention
In this study, we explore the integration of the Retentive Network model, originally designed for Natural Language Processing, into the realm of EEG signal denoising. It is important to emphasize that while the Retentive Network has demonstrated remarkable success in natural language processing, our investigation centers on its applicability to EEG denoising. The Retentive Network is comprised of L identical modules arranged in a stacked fashion, featuring residual connectivity and pre-LayerNorm akin to the Transformer architecture. Each Retnet module comprises two sub-modules: the Multi-scale Retention (MSR) module and the Feedforward Network (FFN) module. For a given input sequence $s=s_{1}s_{2}s_{3}\dots s_{l_{s}}$ (where $l_{s}$ denotes the length of the sequence), the input vector is initially transformed to $X^{0}=\left[x_{1},x_{2},\dots,x_{l_{x}}\right]\in\mathbb{R}^{l_{x}\times d_{model}}$, where $d_{model}$ is the hidden dimension. Subsequently, the Retnet Block can be computed for each layer, denoted as $X^{l}=Retnet(X^{l-1}),l\in[1,L]$. The Simple Retention layer is defined as follows [30]:
@@
where $\bar{\Theta}$ is the complex conjugate of $\Theta$ [31] [28], $W_{Q},W_{K},W_{V}\in\mathbb{R}^{l_{x}\times l_{x}}$, $D\in\mathbb{R}^{l_{x}\times l_{x}}$ combines causal masking and exponential decay of relative distances into one matrix.
To achieve a multichannel-like effect, input sequences can be projected to lower dimensions $d$ times, akin to the multiple-header mechanism in Transformer. This method is employed in each Retention layer with multiple headers $h=\frac{d_{model}}{d}$, where $d$ represents the length of the sequences in each header. Each header utilizes distinct $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d\times d}$, constituting the Muti-Scale Retention (MSR) block. Different $\gamma$ hyperparameters are assigned to various heads in MSR, maintaining simplicity with the same $\gamma$ across different layers. Additionally, a swish [13] gate is incorporated to enhance nonlinear features in each layer. Given an input $X$, the mathematical representation of the Muti-Scale Retention is provided as follows:
@@
Here, $W_{G},W_{O}\in\mathbb{R}^{d_{model}\times d_{model}}$ are learnable parameters, and GroupNorm normalizes the output of each head. Group Normalization (GN) [39] is an alternative to traditional Batch Normalization (BN) [12] that addresses the issue of poor performance with small batch sizes. By normalizing within groups, GN provides a more robust estimation of statistics and helps mitigate the negative impact of small batch sizes or imbalanced data distribution.
@@ 3.3 Signal Embedding
In our extended investigation, it was observed that when the input sequence $s=s_{1}s_{2}s_{3}\dots s_{l_{s}}\in\mathbb{R}^{l_{s}\times 1}$ is relatively short, direct embedding is feasible. However, in general scenarios, where the time-series information of the signal is usually lengthy, direct embedding incurs high computational complexity, hindering effective network training. To address this, we propose the introduction of a concept termed "patch", involving the amalgamation of a series of consecutive samples into a single input feature. This concept is inspired by speech signal processing, where a solitary sample may inadequately represent the current word, while a segment of samples offers more semantic expressiveness. It is noteworthy that EEG signals frequently encompass extensive temporal information, and signal embedding intelligently captures this temporal data. By grouping consecutive samples into patches, the network better retains context and temporal relationships in the signal, enhancing denoising effectiveness. This approach aligns with speech signal processing, where context is pivotal for accurate speech comprehension. Consequently, this paper introduces signal embedding, a more efficient process tailored to the characteristics of EEG signals.
The complete signal embedding process is illustrated in Figure 2(c). Assuming a given patch size, the original sequence is divided accordingly, reducing the sequence length to $l_{s}//patch\_size$. Subsequently, each patch undergoes reshaping and linear projection to attain the desired hidden dimension. This process not only mitigates computational complexity but also preserves timing information more effectively. It’s crucial to note that the signal embedding used here does not employ positional encoding. This is because the Retention mechanism already incorporates positional encoding considerations, obviating the need for additional positional encoding. Mathematically, it can be expressed as follows:
${s}^{\prime}=Pathchfiy(s)={s}^{\prime}_{1}{s}^{\prime}_{2}{s}^{\prime}_{3}\dots s^{\prime}_{\frac{l_{s}}{p}}\in\mathbb{R}^{\frac{l_{s}}{p}\times p}$ (8)
$X^{0}=Embedding({s}^{\prime};\omega)=[x_{1},x_{2},\dots,x_{|x|}]\in\mathbb{R}^{\frac{l_{s}}{p}\times d_{model}}$ (9)
@@ 4.1 Preliminary
Signals disturbed by noise are acquired through the linear combination of the electrooculogram (EOG) or electromyogram (EMG) with the pristine electroencephalogram (EEG). This procedure can be mathematically represented as Equation (11) [42]. The mixed EEG noise signal is denoted as $y\in\mathbb{R}^{l_{y}}$, where $y$ represents the sequence length. The noise-free EEG signal, denoted as $x\in\mathbb{R}^{l_{x}}$, serves as the ground truth, and $n\in\mathbb{R}^{l_{n}}$ represents ocular artifacts or muscle artifacts. It is important to note that the lengths of each sequence $l_{x},l_{y},l_{n}$ are equal. To control the noise level during mixing, we introduce the hyperparameter $\lambda$, regulating the signal-to-noise ratio (SNR) of the noisy signal. Adjustment of different $\lambda$ values enables effective control of SNR magnitude to adapt to various noise environments. The SNR is calculated using Equation (12), while $\lambda$ is determined by Equation (13), where $RMS(\cdot)$ denotes the root mean square of the sample, $RMS(x)$ is the root mean square of the noiseless EEG signal $x$, and RMS $RMS(\lambda\cdot n)$ is the root mean square of the mixed noise $\lambda\cdot n$. These formulas provide flexible adjustment of the signal-to-noise balance to meet diverse signal quality requirements in specific application scenarios.
@@ 4.2.1 Datasets
To validate our proposed network thoroughly, we utilize a semi-simulated EEG dataset (SS2016) [15], contaminated with ocular artifacts, alongside clean EEG signals. This dataset is notable for containing both clean EEG signals and their contaminated counterparts. Collected from $54$ participants during a closed-eye experiment, the signals are devoid of ocular artifacts. Electroencephalogram (EEG) signals were recorded from $19$ electrodes placed according to the International 10-20 system, with each experiment lasting approximately 30 seconds and sampled at 200 SPS[20]. The dataset, comprising pure and contaminated EEG signals, comprises $54$ matrices, each corresponding to one participant. Each matrix consists of $19$ channels, each representing the signal recorded by an electrode. The number of sampling points for channel signals ranged from 5600 to 8400. Data synthesis involved overlaying EOG data onto EEG signals, resulting in recordings of about $30$ seconds in duration but with varying sample counts. For ease of processing, we segmented the data without overlap into small segments of $512$. Post-segmentation, we obtained 11495 samples, with Figure 3 displaying segments 1-4 for participant $1$.
Subsequently, we calculated the SNR values of each sample, rounding them, and depicted the results in Figure 4. Upon observation, we noted SNR values ranging between -10 to 20, with a relatively large number of samples falling within the -5 to 15 range, and fewer samples at other SNR values, posing challenges for network learning. To address this, we followed the data processing approach of the EEGDenoise dataset, separating the noise (n) from the pure signal (x) and contaminated signal (y). We then multiplied n by different $\lambda$ values to attain noise levels ranging from -7dB to 2dB. However, we observed that when the signal-to-noise ratio of signal pairs in the original dataset was high, such as Figure 3b and Figure 3b, excessively large $\lambda$ values were required to achieve low signal-to-noise levels, resulting in signal amplification beyond realistic levels. Data like this not correspond to actual noisy signals and were deemed unsuitable for training samples. Thus, we discarded pairs of samples with SNR higher than 5dB from the original dataset, resulting in a final sample count of 6716. We assigned different $\lambda$ values to n corresponding to each signal, obtaining noise levels from -7dB to 2dB. This dataset is denoted as the SS2016 EOG dataset. The resulting dataset was partitioned into training and test sets in an 8:2 ratio.
Figure 3: The recorded signals from the first electrode of Participant 1 in SS2016 are displayed in sequence as segments 1 through 4, denoted as (a), (b), (c), and (d) respectively. Each segment comprises 512 samples, with a sampling rate of 200 SPS.
Figure 4: The SNR distribution of the SS2016 dataset after signal segmentation is depicted in the figure. It’s important to highlight that the SNR values are rounded up for statistical simplicity. The horizontal axis represents various SNR levels ranging from -10 to 20, while the vertical axis indicates the number of segments corresponding to each SNR level.
@@
However, it is important to note that the EEGDenoiseNet dataset solely provides raw data, and the data provided by SS2016 is not segmented, necessitating researchers to conduct their own processing. This procedure is relatively intricate, posing inconvenience for the exploration of EEG signal denoising through deep learning methodologies. To address this challenge, this paper undertakes the preprocessing of the dataset and subsequently shares the processed dataset as an open-source resource, now accessible on the Hugging Face Hub. The primary objective of this endeavor is to facilitate researchers access to and utilization of processed data, enabling them to concentrate more on the investigation of EEG deep learning denoising methods without being encumbered by the intricacies of data processing. By providing this dataset as an open-source entity, our aim is to stimulate increased research in EEG signal processing and offer a more streamlined resource for the academic community.
@@ 4.2.2 Train details
Throughout the training phase, the network underwent 5000 epochs to ensure comprehensive learning of dataset features. Furthermore, we configured the batch size to 1000, a standard choice that balances memory utilization and training efficacy. Notably, for accelerated training, we harnessed the computational capabilities of an NVIDIA GeForce RTX 4090 Graphics Processing Unit (GPU). Leveraging the parallel computing power of GPUs significantly augmented the speed of deep learning model training, facilitating rapid experimentation and tuning for researchers.
@@ 4.3.1 Comparing method
    Long Short Term Memory (LSTM) Network:
    A Long Short-Term Memory (LSTM) network, adapted from [11], is considered the benchmark for recurrent neural networks (RNNs) [41]. LSTM is capable of learning long-term dependencies, which aids in distinguishing long-term features in noise and EEG signals. Each EEG sample is sequentially fed into LSTM cells, and the output is derived from the state of each cell through a fully-connected network.
  5. 4.
@@ 4.3.2 Evaluation measures
We assess the denoising outcomes using three methods: Relative Root Mean Squared Error in the temporal domain ($RRMSE_{t}$), $RRMSE$ in the spectral domain ($RRMSE_{s}$), and the correlation coefficient ($CC$) [42]. This selection is grounded in a profound understanding of EEG signal characteristics. Firstly, considering the temporal significance of the EEG signal, we employ $RRMSE_{t}$ to quantify the relative error between the denoised and original signals in the time domain. This method exhibits sensitivity to denoising techniques preserving temporal information. Secondly, as EEG signals encapsulate rich spectral information with research often focusing on specific frequency ranges, $RRMSE_{s}$ is employed to ensure the preservation of features in the frequency domain. Lastly, acknowledging the synergistic activities between different brain regions in EEG signals, $CC$ serves as an evaluation metric. $CC$ reflects the linear relationship between the denoised and original signals, crucial for maintaining vital information about interregional correlation. The combined use of these methods facilitates a comprehensive evaluation of denoising performance across time and frequency domains, considering their adaptability to EEG signal characteristics. The mathematical expressions for $RRMSE_{t}$, $RRMSE_{s}$, and $CC$ are represented in Equation (16), (17), and (18) respectively, where $RMS(\cdot)$ denotes root mean square, $PSD(\cdot)$ denotes power spectral density, and $Cov(\cdot)$, $Var(\cdot)$ represent covariance and variance, respectively.
$RRMSE_{t}=\frac{RMS(F(y)-x)}{RMS(x)}=\frac{RMS(\hat{x}-x)}{RMS(x)}$ (16)
$RRMSE_{s}=\frac{RMS(PSD(F(y))-PSD(x))}{RMS(PSD(x))}=\frac{Cov(\hat{x},x)}{Var(\hat{x})Var(x)}$ (17)
@@ 4.3.3 Ablation study
In this study, we pioneer the application of the Retnet network to EEG signal denoising and introduce the innovative concept of signal embedding. To assess the impact of each hyperparameter on denoising performance, we conduct ablation experiments, marking the inaugural exploration of these techniques. Our initial focus is on the influence of patch size and hidden dimension, investigating their effects on network performance. Table 1 presents quantized results for the network applied to EOG, EMG and SS2016 EOG datasets under various configurations of patch size and hidden dimension hyperparameters. Notably, we observe an enhancement in denoising performance with decreasing patch size while maintaining a consistent hidden dimension. This improvement is attributed to the impact of patched sequence length on the network’s feature extraction capability—smaller patch sizes result in larger mini sequence lengths, preserving more information and thereby improving denoising effectiveness. However, a cautious approach is essential as blindly reducing patch size escalates computational complexity due to increased sequence length. Thus, a delicate balance between denoising performance and computational efficiency is imperative. Furthermore, maintaining the same patch size, an increased hidden dimension corresponds to improved denoising performance, aligning with the intuitive understanding that higher dimensionality facilitates enhanced feature extraction.
Table 1: The effect of patch size and hidden dim on the noise reduction performance of EEGDiR. Note mini sequence length must be $l_{s}//patchsize$, where $l_{s}=512$, with $8$ Heads and $4$ DiRBlock layers.
Patch size Mini seq. length Hidden dim EOG dataset EMG dataset SS2016 EOG dataset
$RRMSE_{t}$ $RRMSE_{s}$ $CC$ $RRMSE_{t}$ $RRMSE_{s}$ $CC$ $RRMSE_{t}$ $RRMSE_{s}$ $CC$
32 16 512 0.339 0.367 0.928 0.556 0.561 0.793 0.357 0.392 0.932
32 16 256 0.353 0.377 0.911 0.569 0.575 0.791 0.397 0.466 0.916
32 16 64 0.382 0.371 0.909 0.572 0.577 0.790 0.420 0.525 0.903
16 32 512 0.327 0.361 0.932 0.532 0.501 0.807 0.315 0.362 0.948
16 32 256 0.348 0.371 0.925 0.598 0.573 0.776 0.357 0.395 0.932
16 32 64 0.374 0.378 0.912 0.654 0.593 0.701 0.401 0.443 0.913
Subsequently, we assess the impact of varying the number of block layers L on network performance. Table 2 displays the denoising quantization results for the model applied to EOG, EMG and SS2016 EOG datasets, with fixed parameters patch size $16$, hidden dimension $512$, and heads $8$. The observations indicate a gradual enhancement in denoising performance with an increasing number of layers. This improvement is ascribed to the benefits of residual connections, whereby a higher number of layers does not lead to overfitting. The increased network depth contributes to superior feature extraction capabilities.
@@
Pathch size EOG dataset EMG dataset SS2016 EOG dataset
$RRMSE_{t}$ $RRMSE_{s}$ $CC$ $RRMSE_{t}$ $RRMSE_{s}$ $CC$ $RRMSE_{t}$ $RRMSE_{s}$ $CC$
4 0.327 0.361 0.932 0.532 0.501 0.807 0.315 0.362 0.948
3 0.356 0.380 0.925 0.578 0.583 0.789 0.381 0.447 0.922
2 0.372 0.383 0.917 0.591 0.596 0.781 0.396 0.461 0.917
1 0.394 0.429 0.908 0.613 0.594 0.766 0.411 0.478 0.911